ebjohnson1:

Skylar needs $56 to buy a new video game. If he doubles the amount of money he has now, he will have enough to buy the game with $8 left over.

WilliamMiller08:

so he needed 56 and has 8 left over if he doubles what he already has; (56+8) / 2 = m 56+8 = 64/2 = 32

CrazyHotGhost:

Skylar is a girl's name.